# Prompt Guidance and Human Proximal Perception for HOT Prediction with Regional Joint Loss

### v1

@@ 4.5 Ablation Study
Various components. In Table 2, all experiments were conducted using only cross-entropy loss to train the network model, with $\alpha$, $\beta$, and $\gamma$ all set to 0. The baseline refers to using only the image encoder and image decoder, where the decoder directly performs convolution and prediction on the features from the last layer of the encoder. +Fine denotes refining features by combining the output of each block from the encoder during upsampling in the image decoder, preventing the loss of low-level texture information. +TE indicates the inclusion of the text encoder module. +HPP represents the addition of the human proximal perception module. It is evident that as more components are added, the performance of the proposed method continues to improve.
@@
Different loss weights. The impact of different loss weights on performance is listed in Table 5. When $\alpha$, $\beta$, and $\gamma$ are all set to 1.0, the performance is significantly reduced. We observed that the early values of $\mathcal{L}^{L}$ and $\mathcal{L}^{G}$ are quite large, causing the model to overlook the original segmentation task and instead focus excessively on the presence of other classes within regions. Consequently, as $\alpha$ and $\beta$ decrease, the model’s performance gradually improves, achieving optimal results at $\alpha=0.3$ and $\beta=0.5$. A similar situation occurs with the BE (binary cross-entropy) loss. If $\gamma$ is too large, the focus of the method shifts away, whereas if $\gamma$ is too small, the image-text accuracy is neglected.
@@ 4.6 Visualization
We also provide qualitative visualizations to evaluate the effectiveness of our method. In Figure 4, DHOT [5] represents a CVPR 2023 method, while PIHOT [27] is the latest state-of-the-art (SOTA) approach, accepted at AAAI 2025. It is evident that our proposed method aligns closely with the ground truth. The discrepancy between DHOT’s results and the actual scene is particularly noticeable. Specifically, when evaluating individuals holding a knife, baseball players, and a woman drinking water (top three rows), DHOT’s predictions significantly deviate from the actual interactions.
Similarly, PIHOT produces poor results in certain cases, such as the second sample in the first row and the first sample in the last row. The primary reason for this issue is that PIHOT fails to enforce category consistency within a region, leading to frequent misclassification errors. In contrast, our framework refines predictions through progressive upsampling and utilizes depth information to locate areas around the human body, suppressing irrelevant regions outside the body. Furthermore, by optimizing with RJLoss, our approach better focuses on contact areas and ensures category consistency across the entire region. For more detailed visualizations, please refer to the supplementary materials.

### v2

@@ 4.5 Ablation Study
Various components. In Table 2, all experiments were conducted using only cross-entropy loss to train the network model, with $\alpha$, $\beta$, and $\gamma$ all set to 0. The baseline refers to using only the image encoder and image decoder, where the decoder directly performs convolution and prediction on the features from the last layer of the encoder. +Fine denotes refining features by combining the output of each block from the encoder during upsampling in the image decoder. TE and DE denote the test encoder and depth model, respectively. “+DM+SAM” denotes the HPP module. The results clearly demonstrate the individual contribution of each component as well as their combined impact.
@@
Different loss weights. The impact of different loss weights on performance is listed in Table 5. When $\alpha$, $\beta$, and $\gamma$ are all set to 1.0, the performance is significantly reduced. We observed that the early values of $\mathcal{L}^{L}$ and $\mathcal{L}^{G}$ are quite large, causing the model to overlook the original segmentation task and instead focus excessively on the presence of other classes within regions. Consequently, as $\alpha$ and $\beta$ decrease, the model’s performance gradually improves, achieving optimal results at $\alpha=0.3$ and $\beta=0.5$.
@@ 4.6 Visualization
In Figure 4, it is evident that our proposed method aligns closely with the ground truth. The discrepancy between DHOT’s results and the actual scene is particularly noticeable. Similarly, PIHOT produces poor results in certain cases, such as the second sample in the first row and the first sample in the last row. In contrast, our framework refines predictions through progressive upsampling and utilizes depth information to locate areas around the human body, suppressing irrelevant regions outside the body. Furthermore, by optimizing with RJLoss, our approach better focuses on contact areas and ensures category consistency across the entire region.